# Project Page: https://yikai-wang.github.io/seele ReS Dataset: https://github.com/Yikai-Wang/ReS Repositioning the Subject within Image

### v2

@@ 3.2 Learning task inversion
i) Generating masks for subject removal: In subject repositioning, the mask for the left void mirrors the subject’s shape, but our goal isn’t to generate the subject within the mask. To create training data for this scenario, for each image, we randomly choose a subject and its mask. Next, we move the mask, as shown by the girl’s mask in the center of Fig. 3(b). This results in an image where the masked region includes random portions unrelated to the mask’s shape. This serves as the target for subject removal, with the mask indicating the original subject location and the ground-truth is background areas.
ii) Generating masks for subject completion: In this phase, SEELE addresses scenarios where the subject is partially obscured, with the goal of effectively completing the subject. To integrate this prior information into the task prompt, we generate training data as follows: for each image, we randomly select a subject and extract its mask. Then, we randomly choose a continuous portion of the mask as the input mask. Since user-specified masks are typically imprecise, we introduce random dilation to include adjacent regions within the mask. As illustrated by the umbrella mask on the right side of Fig. 3(b), such a mask serves as an estimate for the mask used in subject completion.
iii) Learning subject harmonization. In SEELE, we achieve subject harmonization by altering the target of diffusion model. To this end, we take as input the inharmonious image and take as output the harmonious image. Additionally, we replace the unmasked region condition with original inharmonious image. Task prompt mainly influences the cross-attention layers. To adapt the self-attention in the diffusion model to preserve the content of masked region while harmonizing appearance, we introduce LoRA adapters [22]. Our training objective is:
where represents the target harmonized image, and is the input inharmonious image. This allows the diffusion model to gradually harmonize the image during denoising. While we modify the training objective, the generation process remains unchanged. This allows us to still utilize the pre-trained stable diffusion model with the learned harmonize-prompt and LoRA parameters, and seamlessly integrate with other modules. See appendix for details.
@@ 4 Experimental Results and Analysis
Examples of subject repositioning. We present subject repositioning results on real-world images using SEELE in Fig. 5. SEELE works well on diverse scenarios, enabling flexible repositioning subject and direction, and achieves high-fidelity repositioned images.
Competitors and setup on ReS. Google Photos’ Magic Editor isn’t publicly accessible, so we can’t compare it with our method. Since there are currently no publicly available models specifically designed for subject repositioning, we mainly compare with original Stable Diffusion inpainting model (SD). We test SD with different prompts, including i) SD performs unconditional generation; ii)SD uses “inpaint” and “complete the subject”; iii) SD uses “Incorporate visually cohesive and high-fidelity background and texture into the provided image through inpainting” and “Complete the subject by filling in the missing region with visually cohesive and high-fidelity background and texture” for subject removal and completion tasks, respectively. iv) SD uses the LoRA fine-tuning strategy to fine-tune the SD at the same training setup of SEELE. Furthermore, we can incorporate alternative inpainting algorithms in SEELE. Specifically, we incorporate LaMa [61], MAT [38], MAE-FAR [4], and ZITS++ [5] into SEELE. We resize images to 512 pixels minimum for compatibility with standard inpainting algorithms. Note that in this experiment, SEELE does not utilize any pre-processing or post-processing techniques. Standard inpainting algorithms cannot tackle subject repositioning without the incorporation of SEELE.
Qualitative comparison. We present qualitative comparison results in Fig. 6 where a larger version and more results are in the appendix. We add orange subject removal mask and blue subject completion mask in the input image. The SD column is SD guided by simple prompt as this variant performs best. Our qualitative analysis indicates that SEELE exhibits better subject removal capabilities without adding random parts and excels in subject completion. When the moved subject overlaps with the left void, SD fills the void by extending the subject. In contrast, SEELE avoids the influence of the subject, as in the top row of Fig. 6. If the mask isn’t precise, SEELE works better than other methods by reducing the impact of unclear edges and smoothing the area, as in the fourth row. SEELE excels in subject completion than typical inpainting algorithms, as in the second-to-last row. Note that SEELE can be enhanced through the post-processing stage.
@@
Quantitative comparison and user-study. We use Learned Perceptual Image Patch Similarity (LPIPS) as quantitative metric and conduct user-study to evaluate user preference from i) quality: the fidelity of the results; ii) visual-consistency (Consist.): the consistency with surrounding area. Our user study on all ReS dataset involves 100 anonymous surveys, reporting the ratio of top-1 preferred option. Results are in Tab. 1. Compared with other methods, SEELE demonstrates significant enhancements in the quality of manipulated images across all metrics. Particularly for the SD, i) our construction of training mask requires object-level ground-truth segmentation in the training dataset, while public dataset do not have large scale annotated dataset (compared with the LAION dataset [59] used by SD which contains 5B training data.) ii) when the training dataset is limited, the task inversion enjoys superior performance while fine-tuning technique leads to over-fitting and cause worse performance.
| Methods | PSNR | SSIM | FID | LPIPS |
| Co-Mod | 21.09 | 0.84 | 30.04 | 0.17 |
| MAT | 20.68 | 0.84 | 32.44 | 0.16 |
| SD(“NA”) | 20.35 | 0.84 | 29.63 | 0.16 |
| SD(“bkg”) | 20.59 | 0.84 | 29.31 | 0.16 |
| SEELE | 21.98 | 0.87 | 24.40 | 0.13 |
| Methods | SD(“NA”) | SD(“bkg”) | SEELE |
| PSNR | 14.48 | 14.60 | 16.00 |
| SSIM | 0.69 | 0.70 | 0.73 |
| FID | 53.52 | 46.58 | 29.06 |
| LPIPS | 0.35 | 0.34 | 0.31 |
@@
Effectiveness of the proposed task-inversion. To further validate the proposed task-inversion, we conduct experiments on standard inpainting task on Places2 [82] and outpainting task on Flickr-Scenery [10], following the standard training and evaluation principles. Quantitative results is in Tab. 2, showcasing the superiority of the proposed task-inversion on both inpainting and outpainting tasks. We provide details and qualitative results in the appendix.
@@
Influence of different task prompts. We train different task prompts to guide different generation direction. Using wrong prompts for tasks can make the model give bad results. We tested this by comparing results from different learned task prompts. As in Fig. 6(a), using a wrong prompt can change the outcome. For subject removal, remove-prompt can correctly generate with background flowers, while complete-prompt wrongly try to add a fly instead of flowers. For subject completion example of trying to add a bird’s head, remove-prompt only added water, but the complete-prompt added the bird’s head properly. This validate the different generation direction learned by our task prompt.
@@
Ablation of Local Harmonization. To tackle the local harmonization sub-task, we learn the harmony-prompt along with the LoRA parameters. To show the efficacy of each module, we conduct an qualitative ablation study in Fig. 6(b). Naturally, if we disable the LoRA parameters, as we use the inharmonious image as unmasked image condition for the stable diffusion model, the model tends to copy the image without significant modification. If we only use LoRA parameter, it works like the unconditional diffusion model to perform local harmonization, but usually performs over- or under- harmonization. Such a manner works to some extent, but can be enhanced with the learned harmony-prompt.
@@
SEELE w/ X. We assess the effectiveness of various components within SEELE during both pre-processing and post-processing phases. We conduct a qualitative comparison of SEELE’s results with and without the utilization of these components, as in Fig. 8, while a detailed analysis of is provided in the appendix. Failure analysis. As a sophisticated system, the success of SEELE relies on the success of each included module. Particularly, the core challenges of subject repositioning include appearance, geometry, and semantic inconsistency issues. i) SEELE addresses the appearance issue, which encompasses the absence of subjects and shadows, as well as unnatural shadows and boundaries. This is achieved through the innovative methods of subject completion, shadow generation, and local harmonization. ii) To tackle the geometry issue, SEELE employs a depth estimation approach that maintains occlusion relationships and perspective accuracy. iii) For resolving semantic inconsistency, SEELE employs techniques for subject removal and completion. The failure of each specific module may lead to the corresponding inconsistency, and resulting in a less-fidelity image.
@@
Limitations. One significant limitation of SEELE is that when the system performs sub-optimally, manual user intervention becomes necessary to enhance the results. For instance, in cases where segmentation fails, users are required to manually correct the segment mask. Similarly, when the subject is occluded, users must provide a mask of potential regions to complete the subject. The former issue could potentially be mitigated through improvements in the segmentation model. However, the latter challenge necessitates the development of a novel model to address the problem of open-vocabulary amodal mask generation [76]. Currently, there lack available foundation models to support open-vocabulary amodal mask generation. These are potential avenues for future research.
@@ 5 Related Works
Image and video manipulation aims to manipulate images and videos in accordance with user-specified guidance. Among these guidance, natural language guidance, as presented in previous studies [12, 48, 36, 37, 68, 27, 14, 79, 15, 7, 66, 25], stands out as particularly appealing due to its adaptability and user-friendliness. Some research efforts have also explored the use of visual conditions, which can be conceptualized as image-to-image translation tasks. These conditions encompass sketch-based [73, 26, 9, 31, 8, 55, 74], label-based [51, 84, 55, 34], line-based [39], and layout-based [44] conditions. In contrast to image manipulation, video manipulation [30, 70, 16] introduces the additional challenge of ensuring temporal consistency across different frames, necessitating the development of novel temporal architectures [3] . Image manipulation primarily revolves around modifying static images, whereas video manipulation deals with dynamic scenes in which multiple subjects are in motion. In contrast, our paper focuses on subject repositioning, relocating one subject while the rest of the image remains unchanged.
@@
Textual inversion [17] is designed to personalize text-to-image diffusion models according to user-specified concepts. It learns new concepts within the embedding space of text conditions while freezing other modules. Null-text inversion [46] learns distinct embeddings at different noise levels to enhance capacity. Some fine-tuning [57] or adaptation [78, 47] techniques inject visual conditions into text-to-image diffusion models. While these approaches concentrate on image patterns, SEELE focuses on the task instruction to guide diffusion models.
@@
Prompt tuning [35, 43, 42] entails training a model to learn specific tokens as additional inputs to transformer models, thereby enabling model adaptation to a specific domain without fine-tuning the model. This technique been widely used in vision-language models [52, 72, 19]. This inspired us to adapt the text-to-image into task-to-image diffusion model by replacing the text conditions.
@@
Image composition [50] is the process of combining a foreground and background to create a high-quality image. Due to differences in the characteristics of foreground and background elements, inconsistencies can arise in terms of appearance, geometry, or semantics. Appearance inconsistencies encompass unnatural boundaries and lighting disparities. Segmentation [33], matting [69], and blending [77] algorithms can be employed to address boundary concerns, while image harmonization [64] techniques can mitigate lighting discrepancies. Geometry inconsistencies include occlusion and disproportionate scaling, necessitating object completion [76] and object placement [63] methods, respectively. Semantic inconsistencies pertain to unnatural interactions between subjects and backgrounds. While each aspect of image composition has its specific focus, the overarching goal is to produce a high-fidelity image. SEELE enhances harmonization capabilities within a single generative model.

### v3

@@ 3.2 Learning task inversion
i) Generating masks for subject removal: In subject repositioning, the mask for the left void mirrors the subject’s shape, but our goal isn’t to generate the subject within the mask. To create training data for this scenario, for each image, we randomly choose a subject and its mask. Next, we move the mask, as shown by the girl’s mask in the center of Figure LABEL:fig:mask-type. This results in an image where the masked region includes random portions unrelated to the mask’s shape. This serves as the target for subject removal, with the mask indicating the original subject location and the ground-truth is background areas.
ii) Generating masks for subject completion: In this phase, SEELE addresses scenarios where the subject is partially obscured, with the goal of effectively completing the subject. To integrate this prior information into the task prompt, we generate training data as follows: for each image, we randomly select a subject and extract its mask. Then, we randomly choose a continuous portion of the mask as the input mask. Since user-specified masks are typically imprecise, we introduce random dilation to include adjacent regions within the mask. As illustrated by the umbrella mask on the right side of Figure LABEL:fig:mask-type, such a mask serves as an estimate for the mask used in subject completion.
iii) Learning subject harmonization. In SEELE, we achieve subject harmonization by altering the target of diffusion model. To this end, we take as input the inharmonious image and take as output the harmonious image. Additionally, we replace the unmasked region condition with original inharmonious image. Task prompt mainly influences the cross-attention layers. To adapt the self-attention in the diffusion model to preserve the content of masked region while harmonizing appearance, we introduce LoRA adapters (Hu et al., 2021). Our training objective is:
where represents the target harmonized image, and is the input inharmonious image. This allows the diffusion model to gradually harmonize the image during denoising. While we modify the training objective, the generation process remains unchanged. This allows us to still utilize the pre-trained stable diffusion model with the learned harmonize-prompt and LoRA parameters, and seamlessly integrate with other modules.
@@ 4 Experimental Results and Analysis
Examples of subject repositioning We present subject repositioning results on real-world images using SEELE in Figure 4. SEELE works well on diverse scenarios, enabling flexible repositioning, and achieves high-fidelity repositioned images.
Competitors and setup on ReS Google Photos’ Magic Editor isn’t publicly accessible, so we can’t compare it with our method. We mainly compare with original Stable Diffusion inpainting model (SD) v2.0. We test SD with different prompts, including i) SD performs unconditional generation; ii) SD uses "inpaint" and "complete the subject"; iii) SD uses "Incorporate visually cohesive and high-fidelity background and texture into the provided image through inpainting" and "Complete the subject by filling in the missing region with visually cohesive and high-fidelity background and texture" for subject removal and completion tasks, respectively. iv) SD uses the LoRA fine-tuning strategy to fine-tune the SD at the same training setup of SEELE. Furthermore, we can incorporate alternative inpainting algorithms in SEELE. Specifically, we incorporate LaMa (Suvorov et al., 2021), MAT (Li et al., 2022), MAE-FAR (Cao et al., 2022), and ZITS++ (Cao et al., 2023) into SEELE. We resize images to 512 pixels minimum for compatibility with standard inpainting algorithms. Note that in this experiment, SEELE does not utilize any pre-processing or post-processing techniques. Standard inpainting algorithms cannot tackle subject repositioning without the incorporation of SEELE.
Qualitative comparison We present qualitative comparison results in Figure 5 where a larger version and more results are in the appendix. We add orange subject removal mask and blue subject completion mask in the input image. The SD column is SD guided by simple prompt as this variant performs best. Our qualitative analysis indicates that SEELE exhibits better subject removal capabilities without adding random parts and excels in subject completion. When the moved subject overlaps with the left void, SD fills the void by extending the subject. In contrast, SEELE avoids the influence of the subject, as in the top row of Figure 5. If the mask isn’t precise, SEELE works better than other methods by reducing the impact of unclear edges and smoothing the area, as in the fourth row. SEELE excels in subject completion than typical inpainting algorithms, as in the second-to-last row. Note that SEELE can be enhanced through the post-processing stage.
@@
Quantitative comparison and user-study We use Learned Perceptual Image Patch Similarity (LPIPS) as quantitative metric and conduct user-study to evaluate user preference from i) quality: the fidelity of the results; ii) visual-consistency (Consist.): the consistency with surrounding area. Our user study on all ReS dataset involves 100 anonymous surveys, reporting the ratio of top-1 preferred option. Results are in Table 1. Compared with other methods, SEELE demonstrates significant enhancements in the quality of manipulated images across all metrics. Particularly for the SDLoRA, i) our construction of training mask requires object-level ground-truth segmentation in the training dataset, while public dataset do not have large scale annotated dataset (compared with the LAION dataset (Schuhmann et al., 2022) used by SD which contains 5B training data.) ii) when the training dataset is limited, the task inversion enjoys superior performance while fine-tuning technique leads to over-fitting and cause worse performance.
@@
Effectiveness of the proposed task-inversion To further validate the proposed task-inversion, we conduct experiments on standard inpainting task on Places2 (Zhou et al., 2017) and outpainting task on Flickr-Scenery (Cheng et al., 2022), following the standard training and evaluation principles. Quantitative results is in Table 2, showcasing the superiority of the proposed task-inversion on both inpainting and outpainting tasks. We provide details and qualitative results in the appendix.
@@
Influence of different task prompts We train different task prompts to guide different generation direction. Using wrong prompts for tasks can make the model give bad results. We tested this by comparing results from different learned task prompts. As in Figure LABEL:fig:inpaint-vs-complete, using a wrong prompt can change the outcome. For subject removal, remove-prompt can correctly generate with background flowers, while complete-prompt wrongly try to add a fly instead of flowers. For subject completion example of trying to add a bird’s head, remove-prompt only added water, but the complete-prompt added the bird’s head properly. This validate the different generation direction learned by our task prompt.
One might also want to use the LoRA technique to adapt the diffusion model for subject removal and completion. We run the experiments and compare this variant with only using the task inversion prompts. As shown in Tab. 3 (b), the LoRA-fine-tuned variant performs poorly. This shows that: (1) The subject removal and completion tasks are similar to generalized inpainting tasks, meaning that simply training task-specific prompts can effectively guide the diffusion model for these sub-tasks. (2) With the same training setup (using COCO), LoRA fine-tuning may reduce the U-Net’s generalization ability. In contrast, our SEELE method keeps the U-Net frozen, maintaining its generalization ability.
@@
Ablation of Local Harmonization To tackle the local harmonization sub-task, we learn the harmony-prompt along with the LoRA parameters. To show the efficacy of each module, we conduct an qualitative ablation study in Figure LABEL:fig:harmony-ablation. Naturally, if we disable the LoRA parameters, as we use the inharmonious image as unmasked image condition for the stable diffusion model, the model tends to copy the image without significant modification. If we only use LoRA parameter, it works like the unconditional diffusion model to perform local harmonization, but usually performs over- or under- harmonization. Such a manner works to some extent, but can be enhanced with the learned harmony-prompt.
Another choice for local harmonization is using specific local harmonization models. However, the benchmark dataset iHarmony4 (Cong et al., 2020) is usually used to train and test on a image size of , which is smaller than the standarad working resolution in SD 2.0 of size . Furthermore, the local harmonization models trained on smaller image sizes cannot generalize well on larger images. For instance, when we tested one of the SOTA models, DucoNet Tan et al. (2023), trained on iHarmony4, it didn’t work as well as our model in our setup, as shown in Table 3(a). Hence we train image harmonization as part of our own framework. Since our framework is flexible, we can easily switch to a better harmonization model in the future if needed. This is also a benefit of our framework.
SEELE w/ X We assess the effectiveness of various components within SEELE during both pre-processing and post-processing phases. We conduct a qualitative comparison of SEELE’s results with and without the utilization of these components, as in Figure 7, while a detailed analysis of is provided in the appendix.
@@
Failure analysis As a sophisticated system, the success of SEELE relies on the success of each included module. Particularly, the core challenges of subject repositioning include appearance, geometry, and semantic inconsistency issues, as shown in Figure 8. i) SEELE addresses the appearance issue, which encompasses the absence of subjects and shadows, as well as unnatural shadows and boundaries. This is achieved through the innovative methods of subject completion, shadow generation, and local harmonization. ii) To tackle the geometry issue, SEELE employs a depth estimation approach that maintains occlusion relationships and perspective accuracy. iii) For resolving semantic inconsistency, SEELE employs techniques for subject removal and completion. The failure of each specific module may lead to the corresponding inconsistency, and resulting in a less-fidelity image.
@@
Limitations One significant limitation of SEELE is that when the system performs sub-optimally, manual user intervention becomes necessary to enhance the results. For instance, in cases where segmentation fails, users are required to manually correct the segment mask. Similarly, when the subject is occluded, users must provide a mask of potential regions to complete the subject. The former issue could potentially be mitigated through improvements in the segmentation model. However, the latter challenge necessitates the development of a novel model to address the problem of open-vocabulary amodal mask generation (Zhan et al., 2020). Currently, there lack available foundation models to support open-vocabulary amodal mask generation. These are potential avenues for future research.
@@ 5 Related Works
Image and video manipulation aims to manipulate images and videos in accordance with user-specified guidance. Among these guidance, natural language guidance, as presented in previous studies (Dong et al., 2017; Nam et al., 2018; Li et al., 2020a; b; Xia et al., 2021; Karras et al., 2019; El-Nouby et al., 2019; Zhang et al., 2021; Fu et al., 2020; Chen et al., 2018; Wang et al., 2018; Jiang et al., 2021), stands out as particularly appealing due to its adaptability and user-friendliness. Some research efforts have also explored the use of visual conditions, which can be conceptualized as image-to-image translation tasks. These conditions encompass sketch-based (Yu et al., 2019; Jo & Park, 2019; Chen et al., 2020; Kim et al., 2020; Chen et al., 2021; Richardson et al., 2021; Zeng et al., 2022), label-based (Park et al., 2019; Zhu et al., 2020; Richardson et al., 2021; Lee et al., 2020), line-based (Li et al., 2019), and layout-based (Liu et al., 2019) conditions. In contrast to image manipulation, video manipulation (Kim et al., 2019; Xu et al., 2019; Fu et al., 2022) introduces the additional challenge of ensuring temporal consistency across different frames, necessitating the development of novel temporal architectures (Bar-Tal et al., 2022) . Image manipulation primarily revolves around modifying static images, whereas video manipulation deals with dynamic scenes in which multiple subjects are in motion. In contrast, our paper focuses on subject repositioning, relocating one subject while the rest of the image remains unchanged.
@@
Textual inversion (Gal et al., 2022) is designed to personalize text-to-image diffusion models according to user-specified concepts. It learns new concepts within the embedding space of text conditions while freezing other modules. Null-text inversion (Mokady et al., 2022) learns distinct embeddings at different noise levels to enhance capacity. Some fine-tuning (Ruiz et al., 2022) or adaptation (Zhang & Agrawala, 2023; Mou et al., 2023b) techniques inject visual conditions into text-to-image diffusion models. While these approaches concentrate on image patterns, SEELE focuses on the task instruction to guide diffusion models.
@@
Prompt tuning (Lester et al., 2021; Liu et al., 2021b; a) entails training a model to learn specific tokens as additional inputs to transformer models, thereby enabling model adaptation to a specific domain without fine-tuning the model. This technique been widely used in vision-language models (Radford et al., 2021; Yao et al., 2021; Ge et al., 2022). This inspired us to adapt the text-to-image into task-to-image diffusion model by replacing the text conditions.
@@
Image composition (Niu et al., 2021) is the process of combining a foreground and background to create a high-quality image. Due to differences in the characteristics of foreground and background elements, inconsistencies can arise in terms of appearance, geometry, or semantics. Appearance inconsistencies encompass unnatural boundaries and lighting disparities. Segmentation (Kirillov et al., 2023), matting (Xu et al., 2017), and blending (Zhang et al., 2020) algorithms can be employed to address boundary concerns, while image harmonization (Tsai et al., 2017) techniques can mitigate lighting discrepancies. Geometry inconsistencies include occlusion and disproportionate scaling, necessitating object completion (Zhan et al., 2020) and object placement (Tripathi et al., 2019) methods, respectively. Semantic inconsistencies pertain to unnatural interactions between subjects and backgrounds. While each aspect of image composition has its specific focus, the overarching goal is to produce a high-fidelity image. SEELE enhances harmonization capabilities within a single generative model.
Drag-based manipulation Pan et al. (2023) also performs dynamic manipulation on a static image. It works by selecting a point on the object, then dragging it to a new location to adjust the object’s shape, direction, or pose to match the drag. To apply this to real images, the process usually involves reversing the image into an initial latent representation Pan et al. (2023) or noise Shi et al. (2024); Mou et al. (2023a); Luo et al. (2024); Mou et al. (2024); Liu et al. (2024). Features are extracted from this representation to recreate the image, and then manipulated to follow the drag direction. The manipulation is guided by methods like motion supervision Pan et al. (2023), explicit feature replacement Shi et al. (2024), or implicit gradient guidance Mou et al. (2023a); Luo et al. (2024); Liu et al. (2024). Fine-tuning is also used in some approaches to preserve the identity of the original image Shi et al. (2024). The key differences between drag-based manipulation and subject repositioning are: (1) Drag-based manipulation focuses on changing the shape of an object but does not handle subject completion, which is essential for subject repositioning. (2) Inversion-based methods struggle to preserve unchanged areas and the repositioned subject, while our approach regenerates only the necessary regions.
@@ 6 Conclusion
Broader Impact Statement
Our proposed SEELE system aims to address the issue of subject repositioning within single images and will be responsive to user intentions. However, there is a risk that it could be misused to create prank images with malicious intent towards individuals, entities, or objects. To mitigate this, we add watermarks to images generated by our SEELE system to indicate their artificial nature.
Acknowledgments
This work was supported in part by NSFC under Grant (No. 62076067). The computations in this research were performed using the CFFF platform of Fudan University.
Appendix A Appendix
A.1 Additional Examples
A.2 Experimental Setting
SEELE is built upon the text-guided inpainting model fine-tuned from SD 2.0, employing the task inversion technique to learn each task prompt with 50 learnable tokens, initialized with text descriptions from the task instructions. For each task, we utilize the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of , weight decay of , and a batch size of 32. Training is conducted on two A6000 GPUs over 9,000 steps, selecting the best checkpoints based on the held-out validation set.
When addressing subject moving and completion, we employ the MSCOCO dataset (Lin et al., 2014), which provides object masks. For image harmonization, the iHarmony4 dataset (Cong et al., 2020) is utilized, offering unharmonized-harmonized image pairs along with subject-to-harmonize masks. MSCOCO comprises 80k training images and 40k testing images, while iHarmony4 includes 65k training images and 7k testing images. This diversity ensures robustness in training task prompts, guarding against overfitting on specific images.
Cost analysis The core component of SEELE is the pre-trained stable diffusion inpainting model, boasting 865.93 million parameters within its UNet backbone. To tailor this stable diffusion model for subject repositioning, we incorporate three distinct task prompts, each sized at 501024 and has 0.5 million trainable parameters. For the local harmonization task, we introduce the LoRA adapter, which encompasses 5.12 million trainable parameters. It’s worth noting that these newly added parameters are lightweight and introduce no additional inference latency when compared to the stable diffusion backbone.
A.3 Analysis of X in SEELE
Here we provide the analysis of each component used in SEELE.
i) Depth estimation for occlusion becomes crucial when users wish to move a subject from the foreground to the background. It helps estimate and correct the occluded parts, ensuring that the repositioned subject blends seamlessly into the scene. As illustrated in the first row of Figure 7, this depth estimation plays a pivotal role in repositioning objects like the tower behind leaves or people behind a car. Neglecting the occlusion relationship can result in unnatural-looking repositioned subjects and a significant loss of image fidelity.
ii) Depth estimation for perspective comes into play when users want to resize the subject proportionally during repositioning. If this aspect is overlooked, the subject’s size remains fixed, which may contradict user expectations.
iii) Matting primarily addresses issues arising from imprecise masks provided by SAM, particularly when dealing with subjects with ambiguous boundaries. Precise masking is crucial because inaccuracies can lead to information leaking in the final output. For example, in Figure 7, imprecise masking might encourage the gaps to generate unnatural dog fur.
iv) Shadow generation is handled by reusing the generative model within SEELE. In cases where a subject includes shadows, such as the left part in Figure 7, we approach it as a subject completion task. The shadow itself becomes the subject, and we employ a learned complete-prompt to guide the diffusion model. Conversely, when a subject lacks shadows, we can transform it into a local harmonization task by utilizing SEELE’s harmonization model to generate shadows.
v) Local harmonization addresses the challenge of appearance inconsistency. When the illumination statistics change after subject repositioning, it’s essential to adjust the subject’s appearance while preserving its texture. As depicted in Figure 7, SEELE excels at this local harmonization task, ensuring seamless integration into the new environment.
A.4 Standard Image Inpainting and Outpainting
Image inpainting The proposed task-inversion approach not only specializes the inpainting model for specific tasks but also enhances its standard inpainting capabilities. We substantiate this claim through experiments conducted on the Places2 dataset (Zhou et al., 2017), where we train SEELE using standard inpainting prompts and compare its performance with other inpainting algorithms. The results are presented in Tab. 2(a) in our paper. Additionally, we provide visual representations of the results in Figure 11, demonstrating SEELE’s advantage in reducing hallucinatory artifacts.
Image outpainting Another commonly used manipulation task involves extending the image beyond its original content. This approach shares a similar concept with subject completion, but it takes a more holistic perspective by enhancing the entire image. We have also conducted experiments on the outpainting task and demonstrated the effectiveness of task inversion. Our experiments were carried out using the Flickr-Scenery dataset (Cheng et al., 2022), and the results are compared with stable diffusion in Tab. 2(b) in our paper. The results indicate the superiority of task inversion employed in SEELE. Furthermore, we provide visual examples for qualitative assessment in Figure 12.
A.5 Necessity of Using Different Datasets to Train SEELE
Our training of the SEELE model utilized only two datasets: COCO, which provides ground-truth object segmentation masks, and iHarmony4, which offers paired images for local harmonization tasks. These datasets, chosen for their public availability, aptly fulfill the varying requirements of different generative sub-tasks. Our training approach, which encompasses both subject movement and completion, employs a unified task inversion technique. Given that local harmonization focuses on not introducing new details in masked areas, we have modified the diffusion model to integrate the characteristics of the masked region, ensuring it aligns with the task’s specific needs.
A.6 Integrating LoRA
When the LoRA adapter is trained, we load them along with the frozen stable diffusion model. As LoRA is implemented as additive layers with the original layers. For example, suppose for a particular layer with input and output . The original stable diffusion performs , while LoRA is trained to perform and only learn while freezing . Then we could introduce a scale hyper-parameter for a trained model When SEELE performs the sub-tasks in manipulation process, we set the lora scale as to preserve the original outputs of stable diffusion. While in the local harmonization process, we set the lora scale as to perform local harmonization. In this regard, we could use the same stable diffusion backbone and perform different sub-tasks using different sub-task prompts (and LoRA parameters).